OpenStudy (anonymous):

Find the sum of the arithmetic sequence. 5, 7, 9, 11, ..., 23 59 40 25 140

OpenStudy (anonymous):

7-5=2; 9-7=2;11-9=2 so 11+2=x; x+2=y; y+2=z and so on. :)

OpenStudy (anonymous):

5,7,9,11,13,15,17,19,21,23 check you answers :) what is 59,40,25 and 140 for?

OpenStudy (anonymous):

i dont get it????

OpenStudy (anonymous):

i think its 140

OpenStudy (hunus):

5+23 = 28 7+21 = 28 9 + 19 = 28 ... 13+15 = 28 The equation will be your first term plus your last term multiplied by the amount of terms and divided by two

OpenStudy (anonymous):

3 + (2)(1) + 3 + (2)(2) + . . . + 3 + (2)(10) (3)(10) + 2(1 + 2 + 3 + . . . + 10) 30 + (2)(55) 140

OpenStudy (anonymous):

thx for the medal!